# AdaLRS: Loss-Guided Adaptive Learning Rate Search for Efficient Foundation Model Pretraining

### v2

@@ 3.3 Extending AdaLRS to WSD Scheduler
Since AdaLRS simply applies scaling factors on the learning rates produced by the scheduler and conduct LR search only in the early training stages, we can apply it to many mainstream schedulers seamlessly. Considering the generalizability of the conclusion and experiment efficiency, we use WSD hu2024minicpm scheduler instead of Cosine scheduler in our experiments, to demonstrate the compatibility of AdaLRS with other schedulers. We use a linear decay function for the final $10\%$ training steps in the implementation of the WSD scheduler. Experiment results are shown in Figure 3. Similar to the experiment results on cosine learning rate schedulers shown in Figure 2(a)(b)(c), AdaLRS adjusts unreasonable LR settings effectively, and shows desired stability for appropriate configurations.
@@ 4 Related Work
Transferring Hyperparameter across Model Sizes. Tensor Program series work studies the transferability of hyperparameter settings across varying model sizes. Recently proposed $\mu$yang2021tensor and $\mu$Transfer yang2022tensor prove the transferability of hyperparameter settings (including learning rate) across Multi-Layer Perceptron popescu2009multilayer (MLP) and Transformers vaswani2017attention ; devlin2018bert ; brown2020language . Powered by $\mu$Transfer, optimal learning rate search can be performed at proxy models and then transferred to larger ones dey2023cerebras ; hu2024minicpm , reducing the search cost by a large margin. However, searching for the optimal learning rate on proxy models can still be resource-consuming, and the optimal learning rate may shift across model sizes in complicated model designs.
@@ 5 Conclusions
In this paper, we propose AdaLRS, which conducts online optimal learning rate search by optimizing the loss descent velocity. We provide pilot study experiments to validate our approach, as well as theoretical proof for the compatibility of the AdaLRS algorithm. We conduct experiments in a series of foundation model pretraining tasks, demonstrating that AdaLRS effectively adjusts inappropriate learning rates to the vicinity of the optimum and accelerates model convergence speed by a large margin. Starting from excessively small learning rates, AdaLRS achieves comparable and even superior performance with baseline models trained with near-optimal learning rates. AdaLRS is validated to improve model performance with varying model sizes, training paradigms, and base schedulers, exhibiting promising potential to be applied in unexplored pretraining tasks. We refer to Appendix C for the discussion of the limitations.
@@ Appendix A AdaLRS Algorithm
In this section, we show the workflow of the proposed AdaLRS algorithm in detail. As shown in Algorithm 1, AdaLRS tracks the velocity of loss descent in pretraining tasks and computes the velocity as the loss curve slope via the least squares method [7]. Learning rate upscaling is triggered when the velocity decays, followed by several validation steps to measure the loss descent velocity after LR adjustment. We then compare the loss descent velocity after LR adjustment with that corresponding to similar loss values before upscaling. If the velocity becomes larger, the upscaling adjustment is retained; otherwise, the model and optimizer states are restored to the exact step before LR upscaling, followed by a LR downscaling adjustment instead. Restoring training states is designed to stabilize the training process, where the upscaled learning rate may disrupt model parameter distribution during training. After several loops of learning rate adjustment, the resulting learning rate eventually falls in the neighborhood of the optimum.
Input: Initial learning rate $\eta_{0}$, Upscale factor $\alpha>1$, Downscale factor $\beta>1$ ($\gcd(\alpha,\beta)=1$), Scale decay factor $\lambda<1$, Loss slope inspection window size $k$, Loss slope variation threshold $10<\theta<$ 1, Loss slope estimation error $e$, AdaLRS start and end step $t_{\text{start}},t_{\text{end}}$
@@
                         $\theta\leftarrow\frac{\theta+1}{2}$
@@
                   $\mathcal{H},\theta\leftarrow\emptyset,\theta_{0}$
                   Restore model states to step $t+k$
@@ Appendix B Proof for the Validity of LR Scaling Factor Values
In this section, we demonstrate the validity of the design of co-prime LR scaling factors.
@@
Let $\alpha,\beta>1$ be co-prime positive integers, for any positive rational number $r$ and $\epsilon>0$, there exist integers $m,n\geq 0$, so that:
@@
We start by pointing out that $\frac{\ln\alpha}{\ln\beta}$ is irrational. This is rather straightforward because if $\frac{\ln\alpha}{\ln\beta}$ is rational, then $\alpha^{q}=\beta^{p}$, where $p,q$ are positive integers. Since $\alpha$ and $\beta$ are coprime, their prime factorizations share no common primes. However, $\alpha^{q}$ and $\beta^{p}$ must share the same prime factors to be equal, forcing $\alpha^{q}=\beta^{p}=1$. It contradicts with the initial conditions, and thus $\frac{\ln\alpha}{\ln\beta}$ must be irrational.
@@ Appendix C Limitations
We also point out that the near-optimal learning rates used in Fit LR experiments in Section 3 are obtained from grid search pilot experiments. We search for the optimal LR setting with exponentially larger LR in a given range, such as $[2e-6,4e-6,8e-6,2e-5,...,8e-3]$ for LLM experiments. Although sufficiently precise for our experiments, the resulting approximated search result may introduce slight variance in optimal LR estimation.

### v3

@@ 3.3 Extending AdaLRS to WSD Scheduler
Since AdaLRS simply applies scaling factors on the learning rates produced by the scheduler and conducts LR search only in the early training stages, we can apply it to many mainstream schedulers seamlessly. Considering the generalizability of the conclusion and experiment efficiency, we use WSD hu2024minicpm scheduler instead of Cosine scheduler in our experiments, to demonstrate the compatibility of AdaLRS with other schedulers. We use a linear decay function for the final $10\%$ training steps in the implementation of the WSD scheduler. Experiment results are shown in Figure 3. Similar to the experiment results on cosine learning rate schedulers shown in Figure 2(a)(b)(c), AdaLRS adjusts unreasonable LR settings effectively, and shows desired stability for appropriate configurations.
@@
3.4 Robustness Across Hyperparameter Settings
In this part, we investigate the model performance dynamics with different AdaLRS hyperparameters. We take 2B LLM pretraining as an example and conduct experiments with small and large learning rate settings, respectively. As shown in Table 4, models trained with AdaLRS demonstrate marked advantages in training loss and evaluation PPL compared to baselines, showing strong robustness across varying hyperparameter combinations. We also observe a clear trend that, as the LR scaling and decay factors decrease, the final learning rate adjusted by AdaLRS tends to decrease for small-LR settings and increase for large-LR settings, respectively. In the experiments with $\alpha,\beta,\lambda=2,1.67,0.99$ and $\alpha,\beta,\lambda=2,1.67,0.95$, both settings yield nearly the same final learning rate, despite their different hyperparameters. This is because LR overshooting in the former triggers a downscaling adjustment, whereas the $0.95$ decay factor in the latter prevents such adjustment.
Table 4: Model performance of 2B LLMs trained with different AdaLRS hyperparameters. $\alpha$, $\beta$, and $\lambda$ refer to LR upscaling factor, downscaling factor, and decay factor, respectively. Baselines with AdaLRS disabled are shown with “-” as hyperparameters.
Small LR Large LR
$\alpha/\beta$ - $3/2$ $2/1.67$ $1.5/1.43$ $2/1.67$ $2/1.67$ - $3/2$ $2/1.67$ $1.5/1.43$ $2/1.67$
$\lambda$ - $0.99$ $0.99$ $0.99$ $0.95$ $0.9$ - $0.99$ $0.99$ $0.99$ $0.9$
Final LR $2.0e^{-5}$ $3.6e^{-4}$ $3.1e^{-4}$ $1.5e^{-4}$ $3.1e^{-4}$ $2.2e^{-4}$ $2.0e^{-3}$ $3.8e^{-4}$ $5.2e^{-4}$ $7.2e^{-4}$ $7.9e^{-4}$
Training Loss $3.07$ $2.55$ $2.55$ $2.58$ $2.54$ $2.54$ $5.30$ $4.94$ $5.10$ $5.15$ $5.08$
Validation PPL $21.59$ $12.87$ $12.75$ $13.11$ $12.61$ $12.68$ $201.54$ $140.13$ $163.95$ $172.01$ $159.55$
Test PPL $21.61$ $12.89$ $12.77$ $13.13$ $12.63$ $12.70$ $201.36$ $140.07$ $163.81$ $171.85$ $159.47$
@@ 4 Related Work
[…]
@@ 5 Conclusions
In this paper, we propose AdaLRS, an online optimal learning rate search algorithm which optimizes the loss descent velocity to approximate the optimal learning rate. To validate this algorithm, we provide both theoretical and experimental analyzes to show the convexity of foundation model pretraining loss and its slope, as well as their shared optimum in LR. We also provide convergence and complexity analysis of AdaLRS to show its effectiveness and efficiency. Experiments in a series of foundation model pretraining tasks demonstrates that AdaLRS effectively adjusts inappropriate learning rates to the vicinity of the optimum in a single run and accelerates model convergence speed by a large margin. Starting from excessively small learning rates, AdaLRS achieves comparable and even superior performance with baseline models trained with near-optimal learning rates. AdaLRS is validated to improve model performance with varying model sizes, training paradigms, and base schedulers, exhibiting promising potential to be applied in unexplored pretraining tasks. We refer to Appendix C for the discussion of the limitations.
@@ Appendix A AdaLRS Algorithm
In this section, we show the workflow of the proposed AdaLRS algorithm in detail. As shown in Algorithm 1, AdaLRS tracks the velocity of loss descent in pretraining tasks and computes the velocity as the loss curve slope via the least squares method [bjorck1990least]. Learning rate upscaling is triggered when the velocity decays, followed by several validation steps to measure the loss descent velocity after LR adjustment. We then compare the loss descent velocity after LR adjustment with that corresponding to similar loss values before upscaling. If the velocity becomes larger, the upscaling adjustment is retained; otherwise, the model and optimizer states are restored to the exact step before LR upscaling, followed by a LR downscaling adjustment instead. Restoring training states is designed to stabilize the training process, where the upscaled learning rate may disrupt model parameter distribution during training. After several loops of learning rate adjustment, the resulting learning rate eventually falls in the neighborhood of the optimum.
Input: Initial learning rate $\eta_{0}$, Upscale factor $\alpha>1$, Downscale factor $\beta>1$ ($\gcd(\alpha,\beta)=1$), Scale decay factor $\lambda<1$, Loss slope inspection window size $k$, Loss slope decay threshold $0<\theta<1$, Loss slope estimation error $e$, AdaLRS start and end step $t_{\text{start}},t_{\text{end}}$
@@
            Restore model states to step $t+k$
       $\mathcal{H},\theta\leftarrow\emptyset,\theta_{0}$
@@
       $\mathcal{H},\theta\leftarrow\emptyset,\theta_{0}$ Restore model states to step $t+k$
@@
       $\theta\leftarrow\frac{\theta+1}{2}$
       // Narrow the decay threshold
@@ Appendix B Proof for the Validity of LR Scaling Factor Values
In this section, we demonstrate the validity of the multiplicatively independent design of LR scaling factors.
@@
Let $\alpha,\beta>1$ be multiplicatively independent real numbers (for all integers $m,n$, $\alpha^{m}=\beta^{n}\implies m=n=0$), for any positive rational number $r$ and $\epsilon>0$, there exist integers $m,n\geq 0$, such that:
@@
We start by pointing out that $\frac{\ln\alpha}{\ln\beta}$ is irrational. This is rather straightforward because if $\frac{\ln\alpha}{\ln\beta}$ is rational, then $\alpha^{q}=\beta^{p}$, where $p,q$ are positive integers, which contradicts with the multiplicatively independent design, and thus $\frac{\ln\alpha}{\ln\beta}$ must be irrational.
@@ Appendix C Limitations
We also point out that the near-optimal learning rates used in Fit LR experiments in Section 3 are obtained from grid search pilot experiments. We search for the optimal LR setting with exponentially larger LR in a given range, such as [2e6,4e6,8e6,2e5,,8e3][2e-6,4e-6,8e-6,2e-5,...,8e-3] for LLM experiments. Although sufficiently precise for our experiments, the resulting approximated search result may introduce slight variance in optimal LR estimation.